OpenStudy (anonymous):

how much is √5/8 ? is 5√2/2 the answer?

OpenStudy (phi):

is it \[ \sqrt{\frac{5}{8} }\] ?

OpenStudy (anonymous):

yes

OpenStudy (adamaero):

0.79 =/= 3.5

OpenStudy (anonymous):

\(\huge\sqrt\frac{5}{8}=\sqrt\frac{5}{2*4}=\frac{1}{2}\sqrt\frac{5}{2}\)

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

What is the orignal problem? @Kfsv10

Directrix (directrix):

The original trig problem.

OpenStudy (adamaero):

? It's solved already

OpenStudy (anonymous):

@Directrix the problem is cos = √5/8, and tan <0

Directrix (directrix):

The cosine is positve and the tangent is negative so we are in Quad III, downstairs on the right.

Directrix (directrix):

Tan x is the ratio of y to x so we need to get sin x. We are back to the identity: sin^2x + cos^2x = 1

Directrix (directrix):

sin^2x + cos^2x = 1 sin^2x + (√5/8)^2 = 1 @Kfsv10 Next step?

OpenStudy (anonymous):

make 8 2√2 and it would look like this √5/2√2 and then 10/2?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

b